Thrifty Parking Garage has 40 parking spaces. Each space can be used by a car or a motorcycle (no trucks). One day when all 40 spaces are in use, the owner walks through the garage and looks at each vehicle. He tells his assistant, "There are 152 wheels in this garage right now." How many motorcycles are in the garage?
Answer by edjones(8007)

Let c=number of cars and m=number of motorcycles.
c+m=40
4c+2m=152
m=40-c
4c+2(40-c)=152 substitute 40-c for m.
4c+80-2c=152
2c=72
c=36
m=4
.
Ed